OpenStudy (anonymous):

solve for y= 2x-10y=70

OpenStudy (anonymous):

steps: 1) subtract 2x from both sides 2) divide both sides by -10

OpenStudy (anonymous):

and the answer is

OpenStudy (anonymous):

thats what you have to find: try the above steps its easy

OpenStudy (anonymous):

no it is not

OpenStudy (anonymous):

what so difficuilt

OpenStudy (anonymous):

never took an algebra class

OpenStudy (j2lie):

Why do you divide by 10?

OpenStudy (anonymous):

to get your final is 7

OpenStudy (j2lie):

I know you're trying to get two points on a line right?

OpenStudy (anonymous):

oook step1: subtracting 2x from both sides we -10y=70 - 2x step2: dividing both sides -10 y = -7+2x or y = 2x - 7

OpenStudy (j2lie):

I thought y=-7.

OpenStudy (anonymous):

y = -7 is incorrect correct answer is y = 2x -7

OpenStudy (anonymous):

nope

OpenStudy (j2lie):

If you solve for x it would be 3.5

OpenStudy (j2lie):

you're right because you have to substitute for the y.